The coordinates of four points are given by P0 = [ 2 2 0 ]T, P1 = [ 2 3 0 ]T, P2 = [ 3 3 0 ]T and P3 = [ 3 2 0 ]T. Find the equation of the Bezier curve. Also, find points on the curve for u = 0, 0.25, 0.5, 0.75, and 1.
Answer #1

Po = [2 2 0]T Pi = [2 3 037 P2 = [3.3 JT P3 = [3 2 037 ozych pcu) = 2 Pi Binfu) Bin(W) = C (mi) u C (h, i). n! i!(n-1)! n=3 2-* (51+ (3+0+1] 3. + + en + 27 3 11 69 2.156 2,563 27 + 81 + 27 + bu 64 oto to to Plonas) = (2.156 2:563 g] plo:5). Pole-o.s헤ot 옵P+ 짧B 11 + It 해「bo+ 9Pi+ 2Fph+2#g [(②]+1] + 3 [3]+2()] 해 1 해 +13 31 +3) 2 + 27 +3 +54 Ottoto 6. 182 6u 0. 2. 34 2. S63 P